# Fast fitting of phylogenetic mixed effects models

### v1

@@ Discussion
In this article, we developed a new method for quickly fitting phylogenetic random effects models, akin to the approaches in the phyr (A. Ives et al. 2020), Hmsc (Tikhonov et al. 2024), and MCMCglmm (Hadfield 2010) R-packages. Because our models are developed as part of the gllvm R-package (Niku et al. 2023), the software supports a wide range of response types, including for Bernoulli, (zero-inflated) Poisson and negative-binomial, Tweedie, and beta responses (Korhonen et al. 2021; Niku et al. 2017; Hui et al. 2017).
@@
By utilizing approximate likelihood methods, the models are magnitudes faster to fit than in the Hmsc and MCMCglmm R-packages, which are based on Markov chain Monte Carlo methods. By applying two more approximations, namely a low-rank VA approximation and NNGPs (Finley et al. 2019) computation time of the proposed models is reduced as much as possible. Two sets of simulation studies confirm that the proposed combination of approximations still enjoys accurate estimation of the phylogenetic signal parameter, while significantly speeding up model fitting.
@@
Unlike in models fitted with the Laplace approximation, we here mostly fitted the models using VA. In this article we purposefully limited the flexibility of the approximation to fit phylogenetic random effects models even more quickly. The flexibility of VA is determined by the number of hyperparameters introduces to the likelihood, which are estimated together with the model parameters. Especially for VA, we find that a good optimisation routine, besides decent initial values, is critical for successful and speedy convergence of the models. In our experience, limited-memory optimisation algorithms (Nocedal and Wright 2006) often lead to faster model fitting and convergence, and an improved ability to successfully leverage methods for parallel computation. In future research we hope to explore improved optimisation routines (as in van der Veen 2024), and their ability to more efficiently fit mixed effects models.
@@
Using real data on 215 species of wood-inhabiting fungi (Abrego et al. 2022b), we demonstrated application of the methods. One of the main novelties of the proposed approach for fitting phylogenetic random effects models is the ability to estimate phylogenetic signal separately for each covariate, without the need to introduce an additional set of random effects as in the phyr R-package, and while simultaneously facilitating the estimation of correlation between random effects for different environmental covariates. Abrego et al. (2022a) fitted various models with and without traits to determine if phylogenetic structuring of species responses occurred at multiple spatial scales simultaneously: at the forest site-scale and at the scale of deadwood logs. Abrego et al. (2022a) post-hoc estimated covariate-specific phylogenetic signal, while here we estimated the covaritae-specific phylogenetic signal directly as part of the model, and while fitting the models in minutes, rather than days as in the Hmsc R-package. Our results largely confirm their findings, namely that fungi responses to decay stage, precipitation, connectivity, and reserve area are phylogenetically structured. We here also find species responses to precipitation are phylogenetically structured.
@@
The phylogenetic signal parameter represents the degree to which species environmental responses are phylogenetically structured, and the associated variance of covariates the rate at which the responses evolve. When the phylogenetic signal parameter is one, species responses to an environmental covariate are fully phylogenetically structured, and at zero the responses are independent (Pearse, Davies, and Wolkovich 2023). The reason for the presence of phylogenetic signal, or the lack thereof, cannot be discerned from phylogenetic random effects models directly. When there is phylogenetic signal, this can be due to the lack of (phylogenetically structured) traits in the model, because the measured traits are not phylogentically structured, or because they are structured in a way that disagrees with the structuring in species environmental responses. In contrast, a lack of phylogenetic signal may be because traits in the model sufficiently introduce phylogenetic structuring to the species responses, because there is no ongoing evolution in the trait that is species environmental responses, or because evolution moves so rapidly that closely related species are (apparently) so different that their environmental responses are independent.
@@
An interesting avenue for future research is to extend the model to incorporate both phylogenetic attraction and repulsion. In their current form phylogenetic random effect models only incorporate positive species associations, meaning that more closely related species can occur in more similar environments. This makes sense if closely related species segregate their resource utilization. However, two closely related species that share the same resource would be more likely to occur in different environments, i.e., exhibit a negative association. Phylogenetic repulsion can be modelled in the current framework, by prior to model fitting replacing the phylogenetic correlation matrix by its inverse (A. R. Ives and Helmus 2011). Still, the current model is limited to either attraction or repulsion, and ideally its form should be adjusted to allow the data to drive that decision instead.

### v2

@@ Discussion
In this article, we developed a new method for quickly fitting phylogenetic mixed-effects models, akin to the approaches in the phyr (Ives et al. 2020), Hmsc (Tikhonov et al. 2024), and MCMCglmm (Hadfield 2010) R-packages. Because our models are developed as part of the gllvm R-package (Niku et al. 2023), the software supports a wide range of response types, including for Bernoulli, (zero-inflated) Poisson and negative-binomial, Tweedie, and beta responses (Korhonen et al. 2021; Niku et al. 2017; Hui et al. 2017).
@@
By utilizing approximate likelihood methods, the models are magnitudes faster to fit than in the Hmsc and MCMCglmm R-packages, which are based on Markov chain Monte Carlo methods. By stacking approximations, namely a VA approximation with matrix normal covariance structure, combined with low-rank approximation of one of the matrices, and NNGPs (Finley et al. 2019) for calculating a sparse approximation to the phylogenetic precision matrix, computation time of the proposed models is reduced as much as possible. Two sets of simulation studies confirm that the proposed combination of approximations still enjoys accurate estimation of the phylogenetic signal, while significantly speeding up model fitting.
@@
We purposefully limited the flexibility of the VA approximation to fit phylogenetic random effects models more quickly. The flexibility of VA is determined by the number of hyperparameters introduced to the likelihood, which are estimated jointly with the model parameters. Especially for VA, we find that a good optimisation routine, besides decent initial values, is critical for successful and speedy convergence of the models. In our experience, limited-memory optimisation algorithms (Nocedal and Wright 2006) often lead to faster model fitting and convergence, and an improved ability to successfully leverage methods for parallel computation. In future research we hope to explore improved optimisation routines (as in van der Veen 2024; Kanzow and Steck 2023), and their ability to more efficiently fit mixed-effects models. In our simulation study the VA approximation was competitive in computational speed to the Laplace approximation as applied using Template Model Builder (Kristensen et al. 2015), though this was limited to 200 species and five covariates, while we expect the VA approximation to especially outperform the Laplace approximation with a higher number of covariates and species. When reducing the number of parameters in the VA approximation as much as possible, the method fits faster than when models are fitted using the Laplace approximation.
@@
The cost of quickly fitting the models comes in form of selecting the number of nearest neighbours for the NNGP approximation. The models are sensitive to the number of neighbours, which thus needs to be carefully assess prior to model fitting. Fortunately, this can straightforwardly be done by quantifying the approximation error under different orderings of species in the data, and numbers of nearest neighbours, and by picking the best combination. We demonstrated picking the number of nearest neighbours and species’ ordering using real data on 215 species of wood-inhabiting fungi (Abrego et al. 2022b). One of the main novelties of the proposed approach for fitting phylogenetic random effects models is the ability to estimate phylogenetic signal separately for each environmental covariate, without the need to introduce an additional set of random effects as in the phyr R-package, and while simultaneously facilitating the estimation of correlation between random effects for different environmental covariates. Abrego et al. (2022a) fitted various models with and without trait covariates to determine if phylogenetic structuring of species responses occurred at multiple spatial scales simultaneously: at the forest site-scale and at the scale of deadwood logs. Abrego et al. (2022a) post-hoc estimated covariate-specific phylogenetic signal, while here we estimated the covariate-specific phylogenetic signal directly as part of the model, and while fitting the models in minutes, rather than days as in the Hmsc R-package.
@@
The phylogenetic signal parameter represents the degree to which species environmental responses are phylogenetically structured, and the associated variance of covariates the rate at which the responses evolve. When the phylogenetic signal parameter is one, species responses to an environmental covariate are fully phylogenetically structured, and at zero the responses are independent (Pearse, Davies, and Wolkovich 2023). The reason for the presence of phylogenetic signal, or the lack thereof, cannot be discerned from phylogenetic random effects models directly. When there is phylogenetic signal, this can be due to the lack of (phylogenetically structured) continuous trait covariates in the model, because the measured trait covariates are not phylogentically structured, or because they are structured in a way that disagrees with the structuring in species environmental responses. In contrast, a lack of phylogenetic signal may be because the trait covariates in the model sufficiently introduce phylogenetic structuring to the species responses, because there is no ongoing evolution in the trait that is species environmental responses, or because evolution moves so rapidly that closely related species are (apparently) so different that their environmental responses are independent.
Various avenues for future research exist to extend the model. Here, we assumed the covariance matrix for the covariate effects to be unstructured. There are various scenarios where a structured covariance matrix would me more sensible, such as for phylogenetically structured site intercepts. Sites are not usually randomly distributed in space or time, while structuring the random effects with an autoregressive or spatial autocorrelation function (e.g., as in Brooks et al. 2017) could address such issues. For computational reasons, such an extension would require an additional approximation to the spatial covariance matrix (similar to the methods used in this article). Although technically possible, combining (control of) further approximations with an intuitive interface might prove challenging.
@@
In their current form phylogenetic random effect models only incorporate positive species associations, meaning that more closely related species can only occur in more similar environments. This makes sense if closely related species segregate their resource utilization. However, two closely related species that share the same resource would be more likely to occur in different environments, i.e., exhibit a negative association. Phylogenetic repulsion can be modelled in the current framework, by prior to model fitting replacing the phylogenetic correlation matrix by its inverse (Ives and Helmus 2011). Still, the current model is limited to either attraction or repulsion, and ideally its form should be adjusted to allow the data to drive that decision instead.
Acknowledgements
The authors thank Sara Taskinen for comments on an earlier version of the manuscript.